OpenStudy (anonymous):

guyz need your help in solving system of linear equation using kirchhoff law .... give me some examples..

OpenStudy (anonymous):

bless you.

OpenStudy (anonymous):

Who sneezed?

OpenStudy (anonymous):

kirchoff

OpenStudy (anonymous):

O_o So confused....

OpenStudy (anonymous):

kirchoffs current law At any node (junction) in an electrical circuit, the sum of currents flowing into that node is equal to the sum of currents flowing out of that node, or: The algebraic sum of currents in a network of conductors meeting at a point is zero. kirchoffs voltage law: This law is also called Kirchhoff's second law, Kirchhoff's loop (or mesh) rule, and Kirchhoff's second rule. The principle of conservation of energy implies that The directed sum of the electrical potential differences (voltage) around any closed network is zero, or: More simply, the sum of the emfs in any closed loop is equivalent to the sum of the potential drops in that loop, or: The algebraic sum of the products of the resistances of the conductors and the currents in them in a closed loop is equal to the total emf available in that loop.

OpenStudy (anonymous):

example of KCL -iT+i1+i2+i3=0 therefore it= i1+i2+i3 ex of kvl v1+v2+v3+v4...-vt=0 vt=v1+v2+v3+...

OpenStudy (anonymous):

ok thnx buddy sniffles :D

OpenStudy (anonymous):

thanks dude